Phone and water bills together cost a company $540 for March. If the phone bill costs 4 times as much as the water bill, find each other.
Answer by Maths68(1474)

Let
Water bills=x
Phone bills costs four times as much as the water bill = 4x
Sum of bills = Total Cost
x+4x=540
5x=540
x=540/5
x=108
water bills = 108
Phone bills=Total cost - water bills cost
Phone bills=540 - 108
Phone bills=432